1.) Increase 9900 in the ratio 9:4

2.) A certain amount of money is divided among A,B and C on the ratio 10:4:3. If B and C together receive $665, what amount is divided among the three

3.) At the market you pay $45,50 for half a kilogram of fish.What would you pay for 2,25kg?


1
Expert's answer
2021-06-21T06:12:09-0400

1)


"9900\cdot(\dfrac{9}{4})=\dfrac{9900\cdot9}{4}=\dfrac{89100}{4}=22275"



2)

Let "x=" the amount of money in 1 part. Then A receives "10x" dollars, B receives "4x" dollars, and C receives "3x" dollars.

Given B and C together receive $665


"4x+3x=665"

"7x=665"

"x=95"

Then the total amount is


"(10+4+3)(95)=1615"

1615 dollars are divided among A,B and C.


3)


"\begin{matrix}
 0.5\ kg & & \$45.50 \\
 2.25\ kg & & \$x
\end{matrix}"

"x=\dfrac{2.25(45.50)}{0.5}"

"x=\$204.75"

We would pat $204.75 for 2.25 kg of fish.